OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. John pushes his bike along the sidewalk at a constant speed. If he increases the force on the bike, what happens? A. The bike becomes more massive. B. The bike loses inertia. C. The bike accelerates. D. The bike gains inertia.

OpenStudy (anonymous):

Well, we know that the bike will not gain mass from an external force, so it's not A. Inertia is defined as: "Inertia is the resistance of any physical object to any change in its motion (including a change in direction). In other words, it is the tendency of objects to keep moving in a straight line at constant linear velocity." And since the force is not causing the bike to shift direction, B and D do not apply. Now what's left is C. If we think about Forces, a force is defined as $$Force = mass * acceleration$$ Since the mass of the bike remains the same no matter how hard you push on it, if the force increases, the acceleration must increase. For example, if the bike had a mass of 5kg, and it was moving at an acceleration of 4 m/s^2, the equation would look like: $$F = (5kg)(4 m/s^2) = 20 N$$ So say we now knew the force being applied to the bike was 40 N. Well, we know the mass of the bike (5kg) didn't change, so the only think left to solve for is a: $$40N = (5kg)*a \to \frac{40N}{5kg} = a \to 8 m/s^2 = a$$ Thus, we can see the acceleration increased, which was the factor that increased the force.

OpenStudy (anonymous):

Protip: Seek the explanation, not the answer. It's much more valuable :)

OpenStudy (anonymous):

thanks so much c: @LolGoCryAboutIt

OpenStudy (anonymous):

Not a problem!